A 30-cm tall solid cylinder of diameter 20 cm is s
Physics help
Answer #1

11.) Total volume of the cylinder

=\pi r^2h= \pi \times 0.1^2 \times 0.3m^3=9.425 \times 10^{-3}m^3

If the force exerted on the ground by the weight of the cylinder is 500 N, then mass of the cylider is

=\frac{500}{9.8}kg=51.02kg

Therefore, density of the material of the cylinder

\frac{51.02}{9.425 \times 10^{-3}}kg/m^3=5413.26kg/m^3=5.413gm/cc

12.) Component of the force acting on the sphere perpendicular to the plane = Mg\cos 35^0

Component of the force acting on the sphere parallel to the plane = Mg\sin 35^0

Torque= Moment of inertia x angular acceleration

N=F \cdot R=I \cdot \alpha \\ \\ \therefore \alpha=\frac{F\times R}{I}=\frac{Mg \sin 35^0 \times R}{\frac{2}{5}MR^2}=\frac{5}{2R}g \sin 35^0

Linear acceleration of the sphere

a=\alpha \cdot R=\frac{5}{2}g \sin 35^0

Linear velocity of the sphere when it will reach the ground will be

v^2=u^2+2as \downarrow \\ \\ v=\sqrt{2 \times a \times s}=\sqrt{2 \times \frac{5}{2}g \sin 35^0 \times 7} m/s^2=14.03ms^{-2}